OpenStudy (anonymous):

Find f^-1 for the function f(x)=(1/3)x^3+(5/3)x+2

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

i am going to make a guess that the actual question had to do with finding the derivative of the inverse at a point of course that is just a guess, so i could be wrong

OpenStudy (anonymous):

ah my bad, u're right, it's (f^-1)'

OpenStudy (anonymous):

how'd i guess?

OpenStudy (anonymous):

you do not have to find the inverse for this, so don't even try it what is the number i assume it is not "find \((f^{-1})'(x)\) but rather find \((f^{-1})'(a)\) for some \(a\)

OpenStudy (anonymous):

yes u're right >,< but because I just want to know how to find the inverse?

OpenStudy (anonymous):

like i said, good luck with that program it is not going to work you are not going to solve a cubic equation for \(y\) what number is it? i can show you how to solve the problem the question is not about finding the inverse of the function

OpenStudy (anonymous):

the number is 12

OpenStudy (anonymous):

whew now we can do it!

OpenStudy (anonymous):

\[\left(f^{-1}\right)'(12)=\frac{1}{f'(f^{-1}(12))}\] so you need two numbers, namely \(f^{-1}(12)\) and \(f'(that)\)

OpenStudy (anonymous):

\(f^{-1}(12)\) you find by inspection i.e. guess at a solution to \[\frac{1}{3}x^2+\frac{5}{3}x+2=12\]

OpenStudy (anonymous):

typo there\[\frac{1}{3}x^3+\frac{5}{3}x+2=12\]

OpenStudy (anonymous):

oh damn no, i cannot solve it did i read the function correctly, or did i make a typo?

OpenStudy (anonymous):

the function is correct, it's here http://i.imgur.com/bpWLqsv.png

OpenStudy (anonymous):

yeah i see that i am going to guess there is a mistake somewhere in the question you do not know how to solve \[\frac{1}{3}x^3+\frac{5}{3}x+2=12\]

OpenStudy (anonymous):

usually the problem is cooked up so the answer is obvious, but not in this case you would have to solve a cubic equation